OpenStudy (anonymous):

z/5=6/30 a 1/5 b 1 c 5 d 25

OpenStudy (lolacole12):

Let's solve your equation step-by-step. z 5 = 6 30 Step 1: Cross-multiply. z =15 = 6 30 z*(30)=(6)*(5) 30z=30 Step 2: Divide both sides by 30. 30z/30=30/30 z=1

OpenStudy (lolacole12):

medal please @jayjaysimmer

OpenStudy (chantysquirrel1129**):

*gasp*